OpenStudy (anonymous):

Rotate the axes to eliminate the xy-term in the equation. Then write the equation in standard form. 61x^2 - 22xy + 61y^2 - 1800 = 0

OpenStudy (nurali):

Suppose we need to rotate the ellipse clockwise by the angle α. Rotating by the angle α moves the point (x,y) to the point (x cos(α) − y sin(α), y cos(α) + x sin(α)) Let's plug this into the equation: 61(x cos(α) − y sin(α))² - 22(x cos(α) − y sin(α))(y cos(α) + x sin(α)) + 61(y cos(α) + x sin(α)) - 1800 = 0 When you simplify this equation using algebra and the dust settles, you should get this: (61 - 22cos(α)sin(α))x² + (61 + 22cos(α)sin(α))y² - 22(cos²α - sin²α)xy = 1800 Now, the xy term has to be zero in order for the ellipse to be upright. Therefore, cos²α = sin²α. The simplest solution for this is α = 45 degrees. Plugging this in, we get: 50x² + 72y² = 1800 x²/36 + y²/25 = 1 x² / 6² + y² / 5² = 1

OpenStudy (anonymous):

thanks so much!

OpenStudy (nurali):

My pleasure.

OpenStudy (anonymous):

Can you help me with this one? Rotate the axes to eliminate the xy-term in the equation. Then write the equation in standard form. 97x2 + 130xy + 97y2 - 2592 = 0